Solve by the elimination method.
4x+6y=14
3x-6y= -21

Answer by Kasko(2)

Solve by the elimination method.
4x+6y=14
3x-6y= -21
7x=-7
X=-1
3x-6y=-21
3(-1)-6y=-21
-3-6y=-21
+3. +3
-6y=-18
/-6. /-6
y=3
(-1,3)
CHECK
4(-1)+6(3)=14
14=14
3(-1)-6(3)=-21
-21=-21